# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v2

@@ Abstract.
Hybrid storage systems (HSS) combine multiple storage devices with diverse characteristics to achieve high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which rearranges stored data (i.e., prefetches hot data and evicts cold data) across the devices to sustain high HSS performance. Prior works focus on improving only data placement or only data migration in HSS, which leads to relatively low HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS, and thus significantly improve system performance. We demonstrate the need for multiple reinforcement learning (RL) agents to accomplish our goal. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two lightweight autonomous RL agents, a data-placement agent and a data-migration agent, which adapt their policies for the current workload and HSS configuration, and coordinate with each other to improve overall HSS performance.
[…]
copyright: none
@@ 1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution for high-performance and large-capacity storage requirements of data-intensive applications such as machine learning (DLRMFacebook, ; near_human_3, ; ho2020denoising, ), large language models (meta_llama, ; jiang2024mixtral, ; liu2024deepseek, ), databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). An HSS combines multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). A typical HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
@@
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy determines the best-fit storage device in the HSS for incoming I/O requests. The data-migration policy rearranges data across the storage devices (i.e., prefetches frequently-accessed data to the fast device and evicts cold data to the slow device) to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns.
@@
We identify four key challenges in designing efficient data-place-ment and data-migration policies. First, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently in data-intensive environments, which makes it hard to optimize the policies. Second, the data-placement policy should have a low performance overhead as it operates on the critical path of I/O request handling. Third, the data-migration policy needs to migrate data across storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the two policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
@@
Limitations of prior works. Prior works propose techniques for either data placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data migration (e.g., (doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )), but they provide relatively low performance when employed together or alone. Our motivational study on performance- (cost-) optimized HSS configurations demonstrates that even the state-of-the-art HSS data-placement technique, Sibyl (singh2022sibyl, ), only achieves 49.9% (41.4%) of the performance of an Oracle policy (see §3 for details). We identify two key reasons for such large performance gaps. First, no prior work proposes a holistic approach to optimize both data placement and data migration together. Second, heuristic- and supervised-learning-based policies (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; doudali2019kleio, ; shetti2019machine, ; ren2019archivist, ; lin2011hot, ; lin2017efficient, )) require hand-tuning and may not adapt to changes in workload characteristics and HSS states.
@@
Since there are no prior works that optimize both placement and migration, we combine prior techniques to create four extended data-management techniques (see §3). Our evaluation shows that these extended techniques underperform because they are naive combinations of prior data-placement and data-migration approaches that are oblivious to each other’s decisions. We conclude that for high HSS performance, both data placement and migration policies need to coordinate with each other.
@@
Need for multi-agent RL. We describe the benefits of RL for HSS data management in §3.1. Despite these benefits, a single RL agent cannot effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration leads to 18.8% (27.7%) lower average performance than Sibyl, which performs only data placement, on a performance- (cost-) optimized HSS (see §3.4). This is because learning to optimize data migration interferes with the process of learning an optimal data placement policy as they are two different tasks (See §3.2).
Key Idea. Harmonia employs two coordinating lightweight autonomous RL agents, respectively for data placement and migration, to improve HSS performance. First, we design Harmonia’s data-placement agent by extending the state-of-the-art placement technique, Sibyl (singh2022sibyl, ). To coordinate with the migration agent, we add new state features and improve reward structure for the placement agent. Second, we design Harmonia’s data-migration agent to identify migration candidates and their target devices by continuously monitoring previously-placed pages in HSS. Harmonia assigns rewards to the migration agent based on the impact of its migrations on future HSS performance. The two agents influence each other’s decisions through actions that affect the HSS environment.
[…]
@@
  • We show that prior data-management approaches have low HSS performance because they do not optimize both data placement and data migration, and we quantify this on a real Linux system with various HSS configurations.
@@ 2.2. Reinforcement Learning
1. State is the representation of the environment. Assuming $S$ is the set of all possible states, $s_{t}$in $S$ is the state at time step $t$.
@@ 3.1. Why RL for Data Management in HSS?
RL is a good fit for HSS data management for three key reasons.
First, unlike heuristic and supervised-learning approaches, an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback without human intervention. Prior work Sibyl (singh2022sibyl, ) shows that RL outperforms heuristic-(e.g., (matsui2017design, ; matsui2017tri, )) and supervised-learning-based (e.g.,  (doudali2019kleio, ; shetti2019machine, )) approaches for HSS data placement. However, Sibyl partially uses RL for HSS data-management (i.e., RL for data-placement and a heuristic approach for data migration) that leaves a large gap between its performance and that of an Oracle approach (as we discuss in §3.4).
@@
Third, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort (as partially demonstrated in Sibyl (singh2022sibyl, )).
@@ 3.3. Prior Approaches to HSS Data Management
Our focus in this work is to co-optimize placement and migration in HSS. In this section, we briefly introduce prior HSS data-management techniques. However, there is no existing technique that optimizes both placement and migration in HSS. We discuss different approaches to optimize both placement and migration by combining (or extending) prior HSS data-management techniques.
@@ 3.4. Effectiveness of Prior Techniques
Performance Results. Fig. 3 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare their average request latencies against two ideal approaches, Fast-Only and Oracle. First, in Fast-Only, we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. All values in Fig. 3 are normalized to Fast-Only.
@@
We make five key observations from Fig. 3. First, all prior techniques exhibit significant performance gaps from Oracle for every tested HSS and workload. The best-performing technique, Sibyl, only achieves 49.9% (41.4%) of Oracle’s performance in performance- (cost-) optimized HSS, which strongly implies the need for holistic data management in HSS. Sibyl’s RL-based data-placement technique can adapt to workload and system changes, but its heuristic data-eviction policy on the critical path lowers its performance.
Second, Sibyl outperforms all other prior techniques, improving the average performance over CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%), respectively, in the performance- (cost-) optimized HSS. Sibyl outperforms CDE, K-SVM and RNN-HSS because they use heuristic- or supervised-learning-based approaches, which cannot easily adapt to changes in workload access patterns and HSS conditions.
Third, Sibyl+K-SVM and Sibyl+RNN-HSS underperform compared to Sibyl. A naive combination of prior approaches performs poorly as the placement and migration policies are oblivious to each other’s decisions. This leads to (1) conflicting decisions from the two policies, and (2) delayed data-placement policy convergence.
Fourth, CDE+RL-Migr shows lower performance than Sibyl by 4.6% (15.6%) on average in performance- (cost-) optimized HSS. CDE+RL-Migr shows comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) because the RL-based migration agent proactively prefetches frequently-read data to the fast device, which Sibyl does not do. However, it underperforms in write-intensive workloads (i.e., SYSTOR17) because CDE’s heuristic data-placement policy places most incoming data in the fast storage device without considering device characteristics, resulting in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive placement policy.
Fifth, SAPM shows lower performance compared to Sibyl by 18.8% (27.7%) on average in the performance- (cost-) optimized HSS, even though SAPM uses RL for both data placement and migration. SAPM’s limited performance is due to two key reasons. (1) SAPM’s single RL agent is unable to learn an optimal policy for data placement and data migration concurrently, and (2) SAPM cannot identify migration candidates in a timely manner because it focuses on the data-migration task only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.2.

### v3

@@ Abstract.
Hybrid storage systems (HSS) integrate multiple storage devices with diverse characteristics to deliver high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which dynamically rearranges stored data (i.e., prefetches hot data and evicts cold data) across the devices to sustain high HSS performance. Prior works optimize either data placement or data migration in isolation, which leads to suboptimal HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS, and thus significantly improve system performance. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two lightweight autonomous RL agents, a data-placement agent and a data-migration agent, that adapt their policies for the current workload and HSS configuration while coordinating with each other to improve overall HSS performance.
We evaluate Harmonia on real HSS configurations with up to four heterogeneous storage devices and seventeen data-intensive workloads. On performance-optimized (cost-optimized) HSS with two storage devices, Harmonia outperforms the best-performing prior approach by 49.5% (31.7%) on average. On an HSS with three (four) devices, Harmonia outperforms the best-performing prior work by 37.0% (42.0%) on average. Harmonia’s performance benefits come with low latency (240nsns for inference) and storage overheads (206 KiB in DRAM for both RL agents combined). We will open-source Harmonia’s implementation to aid future research on HSS.
@@ 1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution to meet the high-performance and large-capacity storage requirements of data-intensive applications such as machine learning (DLRMFacebook, ; near_human_3, ; ho2020denoising, ), large language models (meta_llama, ; jiang2024mixtral, ; liu2024deepseek, ), databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). An HSS combines multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). Typically, an HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
@@
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy determines the most suitable (i.e., best-fit) storage device in the HSS for each incoming I/O request. The data-migration policy rearranges data across the storage devices (i.e., prefetches frequently-accessed data to the fast device and evicts cold data to the slow device) to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns and HSS conditions.
@@
We identify four key challenges in designing effective data-placement and data-migration policies. First, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently in data-intensive environments, which makes it hard to optimize the policies. Second, the data-placement policy operates on the critical path of I/O request handling and it should have a low performance overhead. Third, the data-migration policy needs to migrate data across storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the two policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
@@
Limitations of prior works. Prior works propose techniques to optimize either data placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data migration (e.g., (doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) in isolation, which results in poor performance. For example, an optimized data-placement policy cannot sustain long-term performance if the data-migration decisions remain static or oblivious to workload changes. Similarly, an optimized migration policy cannot compensate for sub-optimal placement decisions that cause unnecessary data movement, wasted bandwidth, and increased device wear. Our motivational study on performance- (cost-) optimized HSS configurations demonstrates that even the state-of-the-art HSS data-placement technique, Sibyl (singh2022sibyl, ), only achieves 49.9% (41.4%) of the performance of an Oracle policy (see §3 for details). While Sibyl optimizes its data-placement policy to adapt to workload and system changes, the large performance gap is due to Sibyl’s heuristic data-migration policy that (ii) considers a limited number of workload features (e.g., access frequency of pages) to identify migration data, and cannot adapt to changes in workload characteristics and HSS states, and (ii) performs migration in the critical path of I/O handling, which adds a significant latency overhead.
@@
Since no prior work jointly optimizes both placement and migration, we combine state-of-the-art data-placement and data-migration techniques to create four extended data-management techniques (see §3). Our evaluation shows that these extended techniques underperform compared to their constituent techniques because the policies operate independently, which leads to conflicting decisions, unnecessary data movement, and wasted system resources. We conclude that a naive combination of data placement and migration policies provides sub-optimal performance, and we need to co-design both the policies with coordination to achieve high HSS performance.
@@
Why RL? We choose RL for data management in HSS because it offers four key benefits compared to heuristic and supervised-learning-based approaches. First, unlike heuristic and supervised learning approaches, RL can learn online and dynamically adapt its decision-making for the current workload and HSS conditions, without prior training using labeled data or human intervention. Second, RL can handle complex tasks such as placement and migration, where each action changes the environment. Third, RL is suitable for even random access patterns by learning the performance impact of actions (i.e., placement or migration) on devices rather than directly predicting future I/O requests. Fourth, RL significantly reduces the system architect’s design effort by offering easy extensibility to various workloads and HSS configurations.
Need for multi-agent RL. We describe the benefits of RL for HSS data management in §3.1. Despite these benefits, a single RL agent cannot effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration leads to 18.8% (27.7%) lower average performance than Sibyl, which performs only data placement, on a performance- (cost-) optimized HSS (see §3.4). This is because learning to optimize data migration interferes with the process of learning an optimal data placement policy as they are two different tasks (See §3.2). A multi-agent RL can mitigate this interference by decoupling the tasks, enabling specialized agents to learn policies that achieve higher HSS performance.
Key Idea. Harmonia employs two coordinating lightweight autonomous RL agents, respectively for data placement and migration, to improve HSS performance. First, we design Harmonia’s data-placement agent to identify the most suitable storage device in the HSS for storing incoming pages. It builds upon the state-of-the-art placement technique, Sibyl (singh2022sibyl, ), but extends it to facilitate coordination with the migration agent. For Harmonia’s data-placement agent, we introduce new migration-aware state features and enhance its reward structure to ensure that placement actions effectively utilize the HSS storage devices. Second, we design Harmonia’s data-migration agent to continuously monitor previously-placed pages in HSS and identify migration candidates and determine their target devices. Harmonia assigns rewards to the migration agent based on the impact of its migrations on future HSS performance. The two agents influence each other’s decisions through actions that affect the HSS environment. This decoupled design prevents policy interference, enables each agent to specialize in its task, and improves overall HSS performance.
We evaluate Harmonia on a real Linux system using seventeen data-intensive workloads and various HSS configurations with up to four devices. Compared to the best-performing prior approach, Sibyl (singh2022sibyl, ), on average, Harmonia provides (1) 49.5% (31.7%) performance improvement on a performance- (cost-) optimized HSS, and (2) 37.0% (42.0%) higher performance on an HSS with three (four) devices. Harmonia incurs low latency (240nsns for inference) and storage overheads (206 KiB for both RL agents together).
@@
  • We show that prior data-management approaches have low HSS performance because they do not optimize both data placement and data migration, and we quantitatively demonstrate their limitations on a real Linux system with various HSS configurations.
@@ 2.2. Reinforcement Learning
1. State is the representation of the environment. Assuming $S$ is the set of all possible states, $s_{t}$ $\in$ $S$ is the state at time step $t$.
@@ 3.1. Why RL for Data Management in HSS?
RL is a good fit for HSS data management for four key reasons.
First, unlike heuristic and supervised-learning approaches, an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback without human intervention.
@@
Third, RL is a suitable approach even for random access patterns compared to heuristic and supervised learning because an RL approach learns the expected performance impact of placement or migration on a target storage device instead of predicting future I/O accesses directly. This allows the RL agents to implicitly learn value-based policies, even when the access patterns are random.
To demonstrate the dynamic changes in workload access patterns, we show the virtual addresses accessed and I/O request sizes during a 300-second execution of an example workload, ssd-02 (see §5 for evaluation methodology), in Figs. 3(a) and 3(b) respectively.
Figure 3. Virtual addresses accessed (top) and request sizes (bottom) captured during a 300-second execution of ssd-02 workload. Each virtual address accessed is represented as a green dot in the top figure.
We make two key observations: (1) the application accesses a broad range of virtual addresses ($\times1$$0^{6} - 1$$\times$$10^{9}$) without fixed patterns. (2) the I/O request sizes vary from random accesses (i.e., 4KiB) to large sequential accesses (i.e., 128KiB). We observe such dynamic access pattern variations in all our chosen workloads. HSS system conditions can also change dynamically. For example, the device capacity utilization changes with each write operation, and access latencies vary significantly when multiple applications access a device concurrently. Prior heuristic and supervised learning based approaches (see S3.3) make fixed decisions at infrequent intervals, and cannot adapt efficiently to the dynamic variations in workload access patterns and system conditions (see §3.4 for our motivational results). Prior work Sibyl (singh2022sibyl, ) shows that RL outperforms heuristic-(e.g., (matsui2017design, ; matsui2017tri, )) and supervised-learning-based (e.g.,  (doudali2019kleio, ; shetti2019machine, )) approaches for HSS data placement. However, Sibyl partially uses RL for HSS data-management (i.e., RL for data-placement and a heuristic approach for data migration) that leaves a large gap between its performance and that of an Oracle approach (as we discuss in §3.4).
Fourth, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort (as partially demonstrated in Sibyl (singh2022sibyl, )).
@@ 3.3. Prior Approaches to HSS Data Management
Our focus in this work is to co-optimize placement and migration in HSS. In this section, we briefly introduce prior HSS data-management techniques (see Table 1). However, there is no existing technique that optimizes both placement and migration in HSS. We discuss different approaches to optimize both placement and migration by combining (or extending) prior HSS data-management techniques.
Table 1. Comparison of HSS Data-Management Techniques.
Technique
Data
Placement
Data
Migration
Key Details
CDE (matsui2017design, ) Heuristic Heuristic
Both policies are hand-tuned.
Data eviction during
I/O request handling.
RNN-HSS (doudali2019kleio, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
K-SVM (shetti2019machine, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
Sibyl (singh2022sibyl, ) RL Heuristic
Data eviction during
I/O request handling.
Sibyl (singh2022sibyl, )
+ K-SVM (shetti2019machine, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
Sibyl (singh2022sibyl, )
+ RNN-HSS (doudali2019kleio, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
CDE (matsui2017design, )
+ RL-Migr
Heuristic RL
Hand-tuned data-placement policy.
SAPM RL
Single RL agent for both tasks.
Harmonia RL RL
Two coordinating RL agents.
@@ 3.4. Effectiveness of Prior Techniques
Performance Results. Fig. 4 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare their average request latencies against two ideal approaches, Fast-Only and Oracle. First, Fast-Only is an ideal policy in which we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. Oracle makes data-placement and data-migration decisions while taking the device capacities into account. All values in Fig. 4 are normalized to Fast-Only.
@@
We make six key observations from Fig. 4. First, all prior techniques fall significantly short of Oracle for every HSS and workload. The best-performing technique, Sibyl, achieves only 49.9% (41.4%) of Oracle’s performance in performance- (cost-) optimized HSS, highlighting the need for holistic HSS data management. While Sibyl’s RL-based placement can adapt to workload and system changes, its heuristic eviction on the critical path limits performance.
Second, Sibyl outperforms CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%) in the performance- (cost-) optimized HSS, respectively, due to its adaptability via RL, unlike heuristic- or supervised-learning-based approaches, which cannot easily adapt to workload changes or HSS conditions.
Third, Sibyl+K-SVM and Sibyl+RNN-HSS underperform compared to Sibyl, as naive combinations of uncoordinated placement and migration policies can lead to (1) conflicing decisions, and (2) slow convergence.
Fourth, CDE+RL-Migr shows lower performance than Sibyl on average (4.6% (15.6%) in performance- (cost-) optimized HSS). It has comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) due to proactive prefetching of frequently-read data by its RL-based migration agent but suffers in write-intensive workloads (i.e., SYSTOR17) due to its heuristic placement policy, which results in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive placement policy.
Fifth, SAPM underperforms Sibyl by 18.8% (27.7%) on average in the performance- (cost-) optimized HSS, limited by (1) its single RL agent handling both placement and migration tasks, and (2) inability to identify migration candidates in a timely manner because it focuses on migration only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.2.
[…]